OpenStudy (anonymous):

Solve the equation the square root of the quantity x plus 4 minus 3 equals 1 for the variable. Show each step of your solution process.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

my guess is \[\sqrt{x+4}-3=1\]

OpenStudy (marihelenh):

From the equation @misty1212 provided, you would add 3 to both sides.\[\sqrt{x+4}=4\] Then square both sides. \[x+4=16\] Solve for x.